OpenStudy (anonymous):

A recipe calls for 5 8 c of butter. If a chef wants to increase the recipe by one-third, how much butter should the chef use?

OpenStudy (flcon04):

I can help

OpenStudy (flcon04):

Are there any choices to choose from?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

So \(\large \sf \frac{5}{8}\) is 100% of the called for recipe. The chef wants to increase it by 33%, making the amount put in 133% or 1 1/3 or 4/3. So what we would do is multiply \[\large \sf \frac{5}{8} \times \frac{4}{3}\]

OpenStudy (anonymous):

and that would be 5x/6

OpenStudy (anonymous):

So whats the Answers ?

Zeronknight (zeronknight):

@helpcollegekid313 \(\Large\cal\color{lime}{✧~ Hello!~Welcome~✧ ~To~OpenStudy~✧ }\)